600,000An automobile dealer has 600,000$ invested in compact cars & mid sized cars....The profit on the sale of the compact cars is 1% and the profit on the mid sized cars is 16%....how much did the dealer invest in midsized cars if the overall profit on the total investement is 12%
Answer by checkley77(12569)

.01X+.16(600,000-X)=.12*600,000
.01X+96,000-.16X=72,000
.01X-16.X=72,000-96,000
-.15X=-24,000
X=-24,000/-.15
X=160,000 WAS INVESTED IN COMPACT CARS.
600,000-160,000=440,000 AMOUNT INVESTED IN MID SIZED CARS.
PROOF:
.01*160,000+.16*440,000=.12*600,000
1,600+70,400=72,000
72,000=72,000